OpenStudy (anonymous):

Find the value of x to the nearest tenth. tan 2 degrees= 4/x

OpenStudy (souvik):

is it\[\tan 2^{0}=4/x\] ?

OpenStudy (anonymous):

yes

OpenStudy (souvik):

x=?

OpenStudy (anonymous):

ugh i have no idea

OpenStudy (anonymous):

x=4tan(2)?

OpenStudy (souvik):

okay multiply x with both sides...

OpenStudy (anonymous):

Oh wait would it be tan^-1

OpenStudy (souvik):

no .......

OpenStudy (anonymous):

Hahahaha sorry I have no clue what I'm doing

OpenStudy (anonymous):

Never mind I'll just figure it out

OpenStudy (souvik):

tan (2)*x=4/x*x x tan(2)=4 do you get it?

OpenStudy (anonymous):

yes

OpenStudy (souvik):

good....now divide both sides by tan 2

OpenStudy (anonymous):

x=tan(4/2)?

OpenStudy (souvik):

no...\[\frac{ x*\tan2^0 }{ \tan 2^0 }=\frac{ 4 }{ \tan2^0 }\] or x=4/tan 2

OpenStudy (anonymous):

so is .14 the rounded answer?

OpenStudy (anonymous):

wait never mind i got 114.5